### v1

# Explicit Motion Handling and Interactive Prompting for Video Camouflaged Object Detection

Abstract
Camouflage poses challenges in distinguishing a static target, whereas any movement of the target can break this disguise. Existing video camouflaged object detection (VCOD) approaches take noisy motion estimation as input or model motion implicitly, restricting detection performance in complex dynamic scenes. In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD, dubbed EMIP, which handles motion cues explicitly using a frozen pre-trained optical flow fundamental model. EMIP is characterized by a two-stream architecture for simultaneously conducting camouflaged segmentation and optical flow estimation. Interactions across the dual streams are realized in an interactive prompting way that is inspired by emerging visual prompt learning. Two learnable modules, i.e. the camouflaged feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the both streams. The prompt fed to the motion stream is learned by supervising optical flow in a self-supervised manner. Furthermore, we show that long-term historical information can also be incorporated as a prompt into EMIP and achieve more robust results with temporal consistency. Experimental results demonstrate that our EMIP achieves new state-of-the-art records on popular VCOD benchmarks. The code will be publicly available.
1 Introduction
Camouflaged object detection (COD) aims at detecting and segmenting those hidden objects that exhibit high intrinsic similarity to their backgrounds. The inherent complexity of distinguishing camouflaged objects from their surroundings poses unique challenges when compared to general object detection [?] and salient object detection [?]. Recently, it has attracted interest of many researchers and facilities broad applications, e.g., medical image segmentation [??] and industrial inspection [??]. While significant progress has been made in image-based [???] and video-based [???] COD tasks, there still remains substantial room for development due to intrinsic difficulty of the tasks.
Figure 1: Different strategies of motion handling in VCOD: (a) Directly feed optical flow maps [??]; (b) Learn implicit motion cues and subsequently utilize them for mask decoding [?]; (c) The proposed interactive prompting paradigm handles motion cues explicitly using a pre-trained optical flow model, and simultaneously conduct optical flow estimation and camouflaged object segmentation. The fire/snowflake symbols denote that most of the model parameters are learnable/frozen in the proposed scheme.
Image-based COD approaches identify camouflaged targets using a static image. Typically, these targets exhibit strong visual resemblances to their backgrounds in terms of texture, color, or edges, posing a challenge for detection using appearance or geometric cues alone. Consequently, motion cues have been investigated in previous works [??] for video camouflaged object detection (VCOD) task. Fig. 1 summarizes the strategies adopted in handling motion cues by these previous works. As presented in Fig. 1 (a), off-the-shelf motion estimators (e.g., [??]) are directly employed for generating offline optical flow, serving as motion cues for identifying camouflaged objects. Notably, offline motion estimation, particularly in camouflaged scenarios, poses significant challenges for common optical flow estimators, often leading to noisy and inaccurate output optical flow. Such erroneous input could misguide detectors and thereby hinder overall performance. To address this issue, SLT-Net [?] proposes a distinct approach by implicit motion modeling in an online fashion (Fig. 1 (b)). While the overall motion modeling part is learnable, learning reliable motion of camouflaged objects from limited VCOD data may be intractable compared to using extensive training data for training common optical flow estimators [??]. Besides, due to the implicit nature of SLT-Net, no explicit regularization or evaluation can be adhered to the motion part, further making the learned motion less effective and reliable. Also, we believe that beyond motion, appearance cues are an important factor conducive to detection as well, since objects could be motion-free or still.
In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD (Fig. 1 (c)), dubbed EMIP, which handles motion cues explicitly by freezing the upstream optical flow model mostly. EMIP is characterized by a two-stream architecture for simultaneously addressing camouflaged segmentation and optical flow estimation. Inspired by emerging advances of the visual prompt learning [???], we design an interactive prompting scheme to achieve interactions across the dual streams, as shown in Fig. 1 (c). Two modules, namely the camouflage feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the dual streams. Due to the absence of authentic optical flow for current VCOD datasets, we learn the prompt fed to the motion stream by supervising optical flow in a self-supervised manner. We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors and further improve the accuracy. Benefiting from the above elaborate designs, the proposed EMIP  effectively leverages noise-robust motion to detect and segment those video camouflaged objects. The main contributions are summarized below:
  • We propose a novel framework for VCOD, dubbed EMIP, which handles motion cues explicitly using a frozen pre-trained optical flow fundamental model. EMIP is formulated into a novel two-stream architecture for simultaneously conducting camouflaged segmentation and optical flow estimation.
  • Inspired by visual prompt learning, the interactions across the two streams are realized in an interactive prompting way, and we propose two modules, i.e., the camouflaged feeder and motion collector, to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively.
  • We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors.
  • EMIP together with its long-term variant, achieve new state-of-the-art records and outperform previous models by notable margins ($\sim$17.0%/5.5average improvement on F-measure/S-measure over the previous best model SLT-Net).
Figure 2: Overall architecture of the proposed EMIP, which consists of two separate streams: explicit motion modeling stream (upper) and object segmentation stream (lower). We use GMFlow  [?] as the fundamental model to handle motion cues. With the camouflage feeder and motion collector, segmentation and motion prompts are injected into each task-specific stream to compensate essential information. The fire/snowflake symbols indicate that the model parameters in this part or block are designated as learnable/frozen.
2 Related Work
Image-based COD. Methods in this type aim to discern camouflaged objects from a single RGB image. Early COD methods [??] relied on hand-crafted features to find targets hidden in the background. However, with the advent of deep learning, COD methods have undergone substantial advancement in recent years. Inspired by natural predatory behavior, SINet-V2 [?] and PFNet  [?] employed a coarse-to-fine strategy. They first generated a preliminary location map for camouflaged objects and then refined it for segmentation. To enhance performance, several studies integrated auxiliary task into a joint learning framework. MGL [?] combined classification or boundary detection task with COD. ZoomNet [?] employed a zooming in and out strategy to the original inputs and processed appearance features at three different scales. Jia et al. [?] proposed the SegMaR framework, an iterative refinement approach designed to locate, magnify, and detect camouflaged objects. HitNet [?] elevated low-resolution representations by leveraging high-resolution features in an iterative feedback loop, effectively mitigating challenges such as edge blurring and detail degradation.
Video-based COD. For the VCOD task, motion cues are crucial to camouflaged object detection. Bidau et al. [?] introduced a method by approximating various motion models derived from dense optical flow. Lamdouar et al.[?] introduced a video registration and segmentation network to detect camouflaged objects, employing optical flow and a difference image as inputs. However, the utilization of inaccurate optical flow may result in accumulative errors in mask prediction. To address this challenge, Cheng et al. [?] proposed a two-stage model that implicitly models and leverages motion information. Different from the above approaches, we propose to handle motion cues in an explicit way as aforementioned, and a two-stream architecture is designed to conduct both optical flow estimation and segmentation.
Visual Prompt Learning. Recently, prompt learning has emerged as a new paradigm that has significantly enhanced the performance of natural language processing (NLP) tasks [?]. Besides, the prompting paradigm has been adopted in many computer vision tasks [??]. Work [?] modified transformer layers by introducing memory tokens, constituting a set of learnable embedding vectors. VPT [?] employed a similar strategy by applying learnable embedding vectors to transformer encoders, achieving noteworthy performance across various downstream recognition tasks. Based on this idea, ViPT [?] integrated modality-complementary visual prompts for task-oriented multi-modal tracking. The popular Segment Anything Model (SAM) [?] integrates various visual prompts like points, boxes, or masks to achieve tailored object segmentation and exhibits decent zero-shot generalization. Previous researches have mainly focused on specific tasks like classification, tracking, or segmentation. In this paper, inspired by [?], we treat intermediate features from one stream as a prompt for injecting complementary information to the other, and propose an interactive prompting scheme for the VCOD task.
3 Methodology
In this work, we propose an end-to-end trainable network EMIP to jointly optimize camouflaged object detection and optical flow estimation. We input an adjacent frame pair $(I_{t},I_{t-1})$ of a video to EMIP, and output a binary segmentation mask of the reference frame $I_{t}$, together with optical flow estimation $\boldsymbol{V}$. The overall architecture of our EMIP is illustrated in Fig. 2. Note that, for object segmentation stream, a set of features $\{f_{n}^{i}\in\mathbb{R}^{H/{2^{i+1}}\times W/{2^{i+1}}\times C_{i}},n\in\{t,t-1\},i=1,...,4\}$ with different scales are extracted from a vision transformer backbone (i.e., PVTv2-B5 [?]), the same as SLT-Net [?], for fair comparison. $W$, $H$, and $C$ represent the width, height, and channel number, respectively. Similar to [?], we adopt the top-three features ($f_{t}^{2},f_{t}^{3},f_{t}^{4}$) for appearance modeling, and discard the first-layer feature $f_{t}^{1}$.
3.1 Fundamental Model for Motion Modeling
To achieve more effective integration of motion information for the VCOD task, we select GMFlow [?] as our fundamental model, which is trained on $\sim$50k frame pairs for optical flow estimation. As shown in Fig. 2, GMFlow can be delineated into two integral components, i.e., a CNN encoder and a transformer decoder. It initially employs the CNN encoder to capture low-level features such as edges, colors, and textures. Subsequently, the transformer decoder, composed of a sequence of self- and cross-attention layers, predicts optical flow map $\boldsymbol{V}\in\mathbb{R}^{H\times W\times 2}$ and matching distribution $\boldsymbol{M}\in\mathbb{R}^{H\times W\times H\times W}$ simultaneously. $\boldsymbol{V}$ is derived from $\boldsymbol{M}$ through operations such as pixel-grid sampling. Here, we flatten the first two dimensions of $\boldsymbol{M}$ to obtain $\boldsymbol{G}\in\mathbb{R}^{H\times W\times HW}$, and employ $\boldsymbol{G}$ as the motion prompt for interacting with segmentation features, leveraging its detailed pixel-to-pixel matching information. Additionally, the flow map $\boldsymbol{V}$ is utilized to reconstruct frame $I_{t}$ from $I_{t-1}$, leading to the computation of a self-supervised loss. For more details about our fundamental motion estimation model, we refer readers to GMFlow [?].
3.2 Segmentation-to-Motion Prompt
A good prompt can fully exploit the potential of the fundamental model, the same for our EMIP. Therefore, the segmentation prompt has to determine where and how to embed it in the fundamental model, considering the nature of the optical flow estimation task.
Position of prompt. The selection of the prompt position is guided by two key observations. Firstly, the choice of the fundamental model plays a pivotal role. Each fundamental model is tailored with a specific architecture; in the case of GMFlow within our EMIP framework, its transformer blocks synergistically form a functional unit crafted for calculating the similarity of pixel features between two frames. Thus, preserving the integrity of this architecture becomes paramount for the coherent generation of optical flow. Secondly, the architectural arrangement of neural networks involves a stratification where lower layers are dedicated to extracting basic features such as edges, colors, and textures, each exhibiting distinctive characteristics across various modalities [?]. This stratification has been shown by [?] to be particularly advantageous for task like optical flow estimation, emphasizing the importance of low-level visual correspondences. Leveraging this insight, we position the segmentation prompt at the lower layers of both the segmentation and motion modeling streams, with the direction being from segmentation to motion. Further details are discussed in Section 4.3, where we delve into various prompt position choices and their impact on results.
Camouflage feeder. To better incorporate the segmentation prompt into the motion stream, we design camouflage feeder. As shown in Fig. 2, our camouflage feeder is based on cross attention with a residual connection, which takes the flow input feature $f^{flow}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of query, and the segmentation prompt feature $f^{2}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of key and value. The residual connection is to maintain more query-related cues, and $d$ represents the embedding dimension. This process can be written as:
${\hat{f}}^{flow}_{t}={\overline{f}}^{flow}_{t}+\mathtt{GDFN}({\overline{f}}^{flow}_{t})$ (1)
where ${\overline{f}}^{flow}_{t}$ is formulated as:
${\overline{f}}^{flow}_{t}=f^{flow}_{t}+\mathtt{CA}(\mathbf{Q},\mathbf{K},\mathbf{V}),$ (2)
where $\mathtt{CA}$ is defined as:
$\mathtt{CA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathtt{Softmax}(\mathbf{Q}\mathbf{K}^{\intercal}/\sqrt{d})\mathbf{V},$ (3)
and $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ represent query, key, and value matrices, respectively. $\bf Q/\bf K$ and $\bf V$ are derived by applying layer normalization and 3$\times$3 convolution to the input feature maps $f^{flow}_{t}$ and $f^{2}_{t}$, respectively. Here, we employ a Gated-Dconv Feed-forward Network (GDFN) [?] to suppress the aggregation of noise prompt. The GDFN process is defined as:
$\mathtt{GDFN}(\mathbf{X})=\mathtt{C1}\left[\phi(\mathtt{C3}(\mathtt{C1}(\mathtt{LN}(\mathbf{X}))))\odot\mathtt{C3}(\mathtt{C1}(\mathtt{LN}(\mathbf{X})))\right],$ (4)
where $\mathtt{C1}$ and $\mathtt{C3}$ are 1$\times$1 convolution and 3$\times$3 depth-wise convolution, respectively. $\phi$ is GELU function [?] and $\odot$ denotes element-wise multiplication. Note that, $\mathtt{GDFN}$ comprises two separate paths, one of which is activated with GELU function. Subsequently, an element-wise product is applied between these two paths. Literature [?] has demonstrated that the gating mechanism in GDFN can control information flow and yield better performance compared with the conventional feed-forward network (FN) [?] for feature restoration. For the same purpose, to mitigate the impact of the integrated noise prompts, we employ GDFN to control noise information flow. Finally, the resulting segmentation-prompted motion features $\hat{f}^{flow}_{t}$ with both knowledges of motion and appearance can be well adapted to subsequent motion reconstruction.
Prompt learning. To better learn the motion information, inspired by [?], we employ a self-supervised loss to optimize the motion-to-segmentation prompt learning. Further elaboration is provided in Section 3.4.
3.3 Motion-to-Segmentation Prompt
Motion-to-segmentation prompt formally provides an auxiliary motion flow that is both temporally and spatially coordinated with the segmentation stream. After obtaining the segmentation-prompted motion, we need to collect this knowledge back to the segmentation stream. Thus, we introduce another module named Motion Collector to integrate motion prompts into the segmentation stream. Given its functional equivalence of motion collector with the camouflage feeder, we opt to maintain the structural configuration of the interaction block rather than introducing a redesign. Specifically, upon receiving motion feature $\boldsymbol{G}$ (flattening the first two dimensions of $\boldsymbol{M}$), we commence by applying a 3$\times$3 convolution operation to $\boldsymbol{G}$ to adjust the channel dimensions to align with $f^{2}_{t}$. Then we feed the motion-to-segmentation prompt $\hat{\boldsymbol{G}}=\mathtt{Conv}(\boldsymbol{G})$ and $f^{2}_{t}$ to the motion collector to obtain the prompted appearance feature $f_{t\leftarrow t-1}^{2}$.
Subsequently, following that most VCOD works [??] use neighbor connection decoder (NCD) [?] to predict the segmentation map, we use NCD to decode the appearance features. NCD with camouflage feeder can be regarded as the prediction head for the motion fundamental model, similar to the concepts of some previous prompt learning methods [??] that design an extra prediction head after freezing the backbone parameters. Specifically, the appearance feature $f_{t\leftarrow t-1}^{2}$ prompted by motion combined with $f^{3}_{t}$ and $f^{4}_{t}$ are fed to NCD to obtain the predicted map. The optimization of motion-to-segmentation prompt and prediction head is under the constraint of segmentation loss, and more details will be given in Sec. 3.4.
Discussion. The motivation for our prompt learning strategy, freezing the motion stream and fully tuning the segmentation stream, is from the observed performance degradation of the PVT [?] pre-trained on ImageNet when it is directly applied to the VCOD task. Similar challenges are also encountered with large models like SAM [?], which struggle in camouflage scenes [?]. Currently, there lacks a freezable foundational model that excels on the VCOD task. In contrast, we find that the optical flow model could achieve good generalization in most VCOD scenarios, when most of its parameters are frozen.
3.4 Supervision and Loss Function
To ensure optimization for each component of the model, we separately define loss functions for the motion and segmentation stream, and then use these two losses to optimize the entire model jointly. Considering the lack of ground truth (GT) optical flow in video camouflaged scenarios, we propose a self-supervised strategy to learn the optical flow estimation stream. For the sake of brevity and clarity, we refer to [?] and warp the frame $I_{t-1}$ according to the optical flow estimation $\boldsymbol{V}$ to obtain the reconstructed reference frame for $I_{t}$, which is denoted as $\hat{I}_{t}$, i.e., $\hat{I}_{t}=\mathtt{Warp}(\boldsymbol{V},I_{t-1})$. We supervise the optical flow by computing the distance between $I_{t}$ and $\hat{I}_{t}$. Thus, the flow loss $\mathcal{L}_{\text{flow}}$ can be expressed as:
$\mathcal{L}_{\text{flow}}=\mathtt{SSIM}(I_{t},\hat{I}_{t}),$ (5)
where $\mathtt{SSIM}$ is pixel-wise photometric loss [?], which is commonly used for reflecting image distortion from three aspects: brightness, contrast, and structure.
For the segmentation stream, we employ a hybrid loss function [?], the same as SLT-Net [?]. The hybrid loss $\mathcal{L}_{\text{seg}}$ can be defined as $\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{IoU}}+\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{e-loss}}$, which includes IoU loss, binary cross-entropy loss, and enhanced-alignment loss. Finally, we jointly optimize the motion modeling and object segmentation streams, whose total loss is formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{seg}}+\mathcal{L}_{\text{flow}}.$ (6)
3.5 Long-term Consistency Modeling
Figure 3: Long-term modeling scheme (${}^{\dagger}$).
The short-term model captures correlated information between adjacent frames, but due to the impact of noises and the limitation of motion estimation model, the short-term model may be less robust. To address this, long-term consideration can be further introduced to improve results’ consistency and therefore detection accuracy. To utilize long-term historical information, as well as maintain feasible computational cost, a dynamic memory model is designed to extract a long-term prompt, which formulates historical features into the prompt to mitigate short-term prediction errors.
Fig. 3 illustrates our long-term consistency modeling scheme, termed ${}^{\dagger}$. Specifically, during training, we freeze the short-term model after it converges and then add the long-term memory module for further learning. Given a video sequence {${{I}_{1},...,{I}_{t-1},{I}_{t}},t>1$}, we sequentially input current frame ${I}_{t}$ and preceding frame ${I}_{t-1}$ into the short-term model to obtain the appearance features $f^{2}_{t}$ and the motion-to-segmentation prompt $\hat{\boldsymbol{G}}$ between the two frames. The element-wise sum of $f^{2}_{t}$ and $\hat{\boldsymbol{G}}$ is conducted, and the results are then fed to an encoder consisting of two convolutional layers. The processed information is later stored into the memory pool. Considering the computational burden and minor impact of distant frames, we implement a FIFO (First-In First-Out) memory pool with fixed capacity ${L}$. Note that such a fixed capacity is configurable according to practical applications. Referring to the long-term setting of SLT-Net, we set ${L}$ to 5 in this paper. Since the memory pool stores appearance and motion features for both the current and historical frames, the features extracted from the current frame are then employed to query the memory pool using a space-time memory read block (STM) [?], yielding the long-term prompt ${P}_{L}$. Next, ${P}_{L}$ and $f^{2}_{t}$ are fed to the motion collector (re-use the same motion collector of the short-term period) for interaction, and the output of collector is subsequently fed to NCD for decoding.
We still adopt the hybrid loss $\mathcal{L}_{\text{seg}}$, the same as the short-term version, for training the long-term counterpart. It is worth noting that compared with SLT-Net which defines long-term modeling as a sequence-to-sequence problem (requiring all frames of a clip), our scheme utilizes only historical frames regarding the current frame. Hence, our scheme is theoretically more suitable for real-time and practical applications, where only past information is available.
4 Experiments
4.1 Implementation Details and Datasets
For fair comparisons, following the training strategy with [?], we train our model with a two-stage pipeline: first train the backbone on the static training set of COD10K (3,040 images) [?], and then fine-tune the whole model with the temporal components on the training set of MoCA-Mask (19,313 frames) [?]. Evaluation is performed on two widely recognized VCOD benchmarks: CAD (836 frames) [?] and MoCA-Mask test set (3,626 frames). Five evaluation metrics are adopted to assess model performance, i.e., structure measure ($\mathcal{S}_{\alpha}$[?], weighted F-measure ($F_{\beta}^{w}$[?], mean absolute error ($\mathcal{M}$), and mean value of Dice and IoU. The proposed EMIP is implemented by PyTorch [?] on a single NVIDIA 4090 GPU. During the training stage, the batch size is set to 6, and the maximum training epoch is 60. The entire model is optimized using the Adam optimizer [?] with a cosine annealing strategy. The maximum and minimum learning rates, along with the maximum adjusted iterations, are set to 1e-5, 1e-6, and 20, respectively. All the input images are resized to 352$\times$352 and subjected to data augmentation techniques, including color enhancement and random rotation. For inference, EMIP takes any two consecutive frames as input and outputs the segmentation prediction together with the corresponding optical flow estimation. However, for ${}^{\dagger}$, it requires consecutive-frame pairs to be input sequentially, in order to leverage its long-term property.
[…]
MoCA-Mask CAD
[…]
Image-based SINet [?] CVPR’20 0.598 0.231 0.028 0.276 0.202 0.636 0.346 0.041 0.381 0.283
SINet-V2 [?] TPAMI’22 0.588 0.204 0.031 0.245 0.180 0.653 0.382 0.039 0.413 0.318
DGNet [?] MIR’23 0.581 0.184 0.024 0.222 0.156 0.686 0.416 0.037 0.456 0.340
FEDER [?] CVPR’23 0.560 0.165 0.031 0.194 0.137 0.691 0.444 0.029 0.474 0.375
HitNet [?] AAAI’23 0.623 0.299 0.019 0.318 0.254 0.685 0.463 0.031 0.478 0.373
Video-based RCRNet [?] ICCV’19 0.555 0.138 0.033 0.171 0.116 0.627 0.287 0.048 0.309 0.229
MG [?] ICCV’21 0.530 0.168 0.067 0.181 0.127 0.594 0.336 0.059 0.368 0.268
PNS-Net [?] MICCAI’21 0.544 0.097 0.033 0.121 0.101 0.655 0.325 0.048 0.384 0.290
SLT-Net [?] CVPR’22 0.637 0.304 0.027 0.356 0.271 0.696 0.471 0.031 0.484 0.392
${}^{\dagger}$ [?] CVPR’22 0.631 0.311 0.027 0.360 0.272 0.696 0.481 0.030 0.493 0.401
EMIP (Ours) 0.669 0.374 0.017 0.424 0.326 0.710 0.504 0.029 0.528 0.415
${}^{\dagger}$ (Ours) 0.675 0.381 0.015 0.426 0.333 0.719 0.514 0.028 0.536 0.425
Figure 4: Visual comparisons of our models (EMIP and ${}^{\dagger}$) with eight state-of-the-art methods.
4.2 Comparisons with State-of-the-arts
To demonstrate the effectiveness of our method, we compare our model with 10 cutting-edge methods, comprising five image-based and five video-based. To ensure fair evaluation, predictions from these methods are either directly downloaded from the official website or generated using models retrained with open-source code.
Quantitative results are presented in Table 1. Notable performance gains can be observed of EMIP and ${}^{\dagger}$ over recently proposed image-based COD and VCOD models, like HitNet [?] and SLT-Net [?], across all five evaluation metrics. Several visual samples are shown in Fig. 4. EMIP appears to be more effective at simultaneously conducting motion and segmentation, and incorporating both enhanced features further makes it better for detecting camouflaged objects. Additionally, predictions of consecutive frames are listed in Fig. 6. One can see that incorporating the long-term modeling scheme, namely ${}^{\dagger}$, can further refine results and lead to boosted performance.
Table 2: Ablation results on camouflage feeder (CF) and motion collector (MC) modules of EMIP on MoCA-Mask dataset. The best results are highlighted in bold.
[…]
1 - - 0.627 0.268 0.031 0.340 0.254
2 - 0.657 0.349 0.020 0.394 0.300
3 0.669 0.374 0.017 0.424 0.326
4.3 Ablation Studies
We conduct thorough ablation studies by removing or substituting components from the complete EMIP.
Camouflage feeder and motion collector. We first validate the role of the two prompt integration modules, i.e., camouflage feeder and motion collector. As shown in Table 2, we start with the baseline model (#1), which directly uses the segmentation stream to extract single-frame information for prediction, i.e., removing the motion modeling stream of EMIP. Configuration #2 means using only motion information to prompt the segmentation stream, meanwhile discarding segmentation-to-motion prompt. Model #3 is the full EMIP. The integration of motion information is beneficial for the detection of camouflaged objects in video sequences. By incorporating the segmentation-to-motion prompt through the camouflage feeder into the motion stream, we enhance the robustness of motion cues, further resulting in improved performance for segmentation.
Full-tuning v.s. Freezing. Existing models usually address motion modeling by fully tuning all the parameters. In contrast, our EMIP adopts a prompt learning strategy to conduct motion modeling by freezing the motion fundamental model. From the comparisons in Table 3, where both settings are preloaded with pre-trained weights, the superiority of freezing the motion model over full-tuning is evident. This implies the prompt learning strategy for the motion modeling stream can better exploit its potential on limited VCOD data.
Table 3: Different learning strategies for the motion stream.
[…]
Full-tuning 0.645 0.326 0.018 0.364 0.281
Freezing (Ours) 0.669 0.374 0.017 0.424 0.326
Table 4: Different prompt destinations on the motion stream.
[…]
CNN Layer1 0.644 0.331 0.018 0.368 0.285
CNN Layer2 0.650 0.334 0.021 0.379 0.289
CNN Layer3 0.669 0.374 0.017 0.424 0.326
Prompt destination on the motion stream. The prompt destination should address where to feed the appearance prompt on the motion fundamental model. In this experiment, we utilize $f^{2}_{t}$ as both segmentation-to-motion prompt and motion collector input. Then we use it to prompt different CNN layers, and the evaluation results are presented in Table 4. It can be seen that prompting with the features after CNN Layer3 achieves superior performance. The underlying reason is that distinct-level features need the corresponding matching prompt to enhance level-specific representations.
Prompt source from the segmentation stream. To explore the most suitable appearance prompt to feed to the motion stream, we conduct another experiment with different prompt features, i.e., $f^{2}_{t}$, $f^{3}_{t}$, and $f^{4}_{t}$. The experimental results are presented in Table 5, where it is evident that utilizing appearance features $f^{2}_{t}$ as prompt yields the best performance. One possible explanation for this phenomenon is that slight motion of a moving object is likely to be missed in lower-resolution feature maps, namely $f^{3}_{t}$ and $f^{4}_{t}$, and therefore degrading the efficacy for prompting to the motion stream.
Table 5: Different prompt sources from the segmentation stream.
[…]
$f^{2}_{t}$ 0.669 0.374 0.017 0.424 0.326
$f^{3}_{t}$ 0.654 0.350 0.018 0.389 0.300
$f^{4}_{t}$ 0.624 0.289 0.023 0.331 0.251
Prompt destination on the segmentation stream. We leverage motion information to selectively prompt distinct appearance features or simultaneously prompt all features by introducing additional motion collectors. The results, as presented in Table 6, demonstrate that utilizing motion to prompt feature $f^{2}_{t}$ outperforms other features ($f^{3}_{t}$ or $f^{4}_{t}$) or all features collectively. This implies that semantic features at corresponding hierarchical levels between segmentation and motion tasks can mutually enhance the representation of the original features. Yet, the integration of features from disparate levels, which lack alignment, may compromise the fidelity of the original feature representation.
Table 6: Different prompt destinations on the segmentation stream.
[…]
0.669 0.374 0.017 0.424 0.326
0.642 0.319 0.019 0.359 0.271
0.646 0.325 0.021 0.366 0.283
0.639 0.313 0.020 0.361 0.271
Effectiveness of motion self-supervision. Quantitative and qualitative results are reported to validate the effectiveness of different training strategies for the motion stream. Due to the absence of optical flow ground truth, conducting quantitative analyses for the output optical flow is not feasible. Thus, we evaluate its effectiveness via segmentation/prediction performance. From Table 7, the model with a self-supervised loss achieves better performance, which demonstrates that the prompted motion under self-supervision is beneficial to boosting detection performance.
Fig. 5 visualizes comparisons of our model with different settings in terms of optical flow prediction for camouflaged scenarios. The results show that sometimes the original GMFlow cannot perceive camouflaged moving objects (Fig. 5 (c)), whereas ours that incorporates camouflage information with the prompt learning paradigm can better detect the targets (Fig. 5 (a)). Fig. 5 (d) further shows that the lack of self-supervised loss in our EMIP leads to worse flow prediction. In contrast, full-tuning the GMFlow part (Fig. 5 (b), correspond those results in Table 3) is hardly generalized in new camouflaged scenarios.
Figure 5: Visual comparisons employing different designs: (a) Our self-supervision (default EMIP), (b) Full-tuning, (c) GMFlow, (d) Ours w/o self-supervised loss.
Table 7: EMIP without/with the self-supervised loss.
[…]
w/o self-supervised 0.644 0.321 0.020 0.363 0.277
w/ self-supervised 0.669 0.374 0.017 0.424 0.326
Effectiveness of long-term modeling. As aforementioned in Table 1, by introducing the long-term consistency prompt to the short-term model EMIP, we have further enhanced the performance on all five metrics. Additionally, visual comparisons on the same clip in Fig. 6 show that the incorporation of long-term consistency effectively harnesses temporal consistency, thereby reducing errors in short-term prediction.
Figure 6: Visual comparisons on consecutive video frames.
5 Conclusion
We propose EMIP, an innovative framework for VCOD that explicitly handles motion cues through the utilization of a frozen pre-trained optical flow fundamental model. EMIP adopts a novel two-stream architecture, concurrently addressing camouflaged segmentation and optical flow estimation. The core idea of interaction between these two streams is orchestrated through an interactive prompting mechanism. Experimental results show the paradigm of interactive prompting of EMIP can enhance outputs of both streams, further achieving accurate prediction. In addition, we present an extended version of EMIP, incorporating historical features into the prompt to alleviate short-term prediction errors and enhance overall accuracy. Our contributions not only achieve compelling results on two VCOD benchmarks but also provide fresh insights into addressing the challenging VCOD task. We hope that the proposed framework will serve as a catalyst for inspiring further research in this emerging field.

### v2

@@
 Abstract
-Camouflage poses challenges in distinguishing a static target, whereas any movement of the target can break this disguise.
+Camouflage poses notable challenges in distinguishing a static target, as it usually blends seamlessly with the background.
+However, any movement by the target can disrupt this disguise, making it detectable.
 Existing video camouflaged object detection (VCOD) approaches take noisy motion estimation as input or model motion implicitly, restricting detection performance in complex dynamic scenes.
@@
 Furthermore, we show that long-term historical information can also be incorporated as a prompt into EMIP and achieve more robust results with temporal consistency.
-Experimental results demonstrate that our EMIP achieves new state-of-the-art records on popular VCOD benchmarks.
-The code will be publicly available.
-1 Introduction
+By leveraging promoting techniques based on EMIP, the proposed long-term model EMIP incurs lower training cost with only 8.5M trainable parameters (less than 8% of the total model parameters).
+Experimental results demonstrate that both EMIP and EMIP set new state-of-the-art records on popular VCOD benchmarks.
+Additionally, comparative evaluations against other video segmentation models on a wider range of video segmentation tasks demonstrate the robustness and superior generalization capabilities of EMIP.
+Our code is made publicly available at https://github.com/zhangxin06/EMIP.
+Index Terms:
+Video camouflaged object detection, explicit motion modeling, interactive prompting, semantic segmentation, deep learning.
+I Introduction
 Camouflaged object detection (COD) aims at detecting and segmenting those hidden objects that exhibit high intrinsic similarity to their backgrounds.
-The inherent complexity of distinguishing camouflaged objects from their surroundings poses unique challenges when compared to general object detection [?] and salient object detection [?].
-Recently, it has attracted interest of many researchers and facilities broad applications, e.g., medical image segmentation [??] and industrial inspection [??].
-While significant progress has been made in image-based [???] and video-based [???] COD tasks, there still remains substantial room for development due to intrinsic difficulty of the tasks.
-Figure 1: Different strategies of motion handling in VCOD: (a) Directly feed optical flow maps [??]; (b) Learn implicit motion cues and subsequently utilize them for mask decoding [?]; (c) The proposed interactive prompting paradigm handles motion cues explicitly using a pre-trained optical flow model, and simultaneously conduct optical flow estimation and camouflaged object segmentation.
+The inherent complexity of distinguishing camouflaged objects from their surroundings poses unique challenges when compared to general object detection [1] and salient object detection (SOD) [2].
+Recently, it has attracted interest of many researchers and facilities broad applications, e.g., medical image segmentation [3, 4] and industrial inspection [5, 6].
+While significant progress has been made in image-based [7, 8, 9] and video-based [10, 11, 12] COD tasks, there still remains substantial room for development due to intrinsic difficulty of the tasks.
+Figure 1: Different strategies of motion handling in VCOD: (a) Directly feed optical flow maps [11, 13]; (b) Learn implicit motion cues and subsequently utilize them for mask decoding [10]; (c) The proposed interactive prompting paradigm handles motion cues explicitly using a pre-trained optical flow model, and simultaneously conduct optical flow estimation and camouflaged object segmentation.
 The fire/snowflake symbols denote that most of the model parameters are learnable/frozen in the proposed scheme.
@@
 Typically, these targets exhibit strong visual resemblances to their backgrounds in terms of texture, color, or edges, posing a challenge for detection using appearance or geometric cues alone.
-Consequently, motion cues have been investigated in previous works [??] for video camouflaged object detection (VCOD) task.
+Consequently, motion cues have been investigated in previous works [11, 13] for video camouflaged object detection (VCOD) task.
 Fig.
@@
 As presented in Fig.
-1 (a), off-the-shelf motion estimators (e.g., [??]) are directly employed for generating offline optical flow, serving as motion cues for identifying camouflaged objects.
+1 (a), off-the-shelf motion estimators (e.g., [14, 15]) are directly employed for generating offline optical flow, serving as motion cues for identifying camouflaged objects.
 Notably, offline motion estimation, particularly in camouflaged scenarios, poses significant challenges for common optical flow estimators, often leading to noisy and inaccurate output optical flow.
 Such erroneous input could misguide detectors and thereby hinder overall performance.
-To address this issue, SLT-Net [?] proposes a distinct approach by implicit motion modeling in an online fashion (Fig.
+To address this issue, SLT-Net [10] proposes a distinct approach by implicit motion modeling in an online fashion (Fig.
 1 (b)).
-While the overall motion modeling part is learnable, learning reliable motion of camouflaged objects from limited VCOD data may be intractable compared to using extensive training data for training common optical flow estimators [??].
+While the overall motion modeling part is learnable, learning reliable motion of camouflaged objects from limited VCOD data may be intractable compared to using extensive training data for training common optical flow estimators [14, 16].
 Besides, due to the implicit nature of SLT-Net, no explicit regularization or evaluation can be adhered to the motion part, further making the learned motion less effective and reliable.
@@
 EMIP is characterized by a two-stream architecture for simultaneously addressing camouflaged segmentation and optical flow estimation.
-Inspired by emerging advances of the visual prompt learning [???], we design an interactive prompting scheme to achieve interactions across the dual streams, as shown in Fig.
+Inspired by emerging advances of the visual prompt learning [17, 18, 19], we design an interactive prompting scheme to achieve interactions across the dual streams, as shown in Fig.
 1 (c).
@@
   • Inspired by visual prompt learning, the interactions across the two streams are realized in an interactive prompting way, and we propose two modules, i.e., the camouflaged feeder and motion collector, to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively.
-  • We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors.
-  • EMIP together with its long-term variant, achieve new state-of-the-art records and outperform previous models by notable margins ($\sim$17.0%/5.5average improvement on F-measure/S-measure over the previous best model SLT-Net).
-Figure 2: Overall architecture of the proposed EMIP, which consists of two separate streams: explicit motion modeling stream (upper) and object segmentation stream (lower).
-We use GMFlow  [?] as the fundamental model to handle motion cues.
-With the camouflage feeder and motion collector, segmentation and motion prompts are injected into each task-specific stream to compensate essential information.
-The fire/snowflake symbols indicate that the model parameters in this part or block are designated as learnable/frozen.
-2 Related Work
-Image-based COD.
+  • We also propose a long-term variant of EMIP, dubbed EMIP, by formulating historical features into the prompt to mitigate short-term prediction errors.
+  • EMIP together with its long-term variant EMIP, achieve new state-of-the-art records and outperform previous models by notable margins ($\sim$17.0%/5.5average improvement from EMIP on F-measure/S-measure over the previous cutting-edge model SLT-Net).
+The remainder of the paper is organized as follows: Section II discusses related work on image and video-based COD, closely related video object segmentation, salient object detection, and visual prompt learning.
+Section III describes the proposed models in detail.
+Section IV includes experimental results, comparisons, and analyses.
+Finally, conclusions are drawn in Section V.
+II Related Work
+II-A Image-based COD
 Methods in this type aim to discern camouflaged objects from a single RGB image.
-Early COD methods [??] relied on hand-crafted features to find targets hidden in the background.
+Early COD methods [20, 21] relied on hand-crafted features to find targets hidden in the background.
+Li et al.
+[22] proposed a texture guided weighted voting strategy to detect camouflaged objects.
+Then they further introduced a fusion framework [23] to address camouflaged problems in the wavelet domain.
+And more, Garcia [24] review the background subtraction methods and outlook future direction.
 However, with the advent of deep learning, COD methods have undergone substantial advancement in recent years.
-Inspired by natural predatory behavior, SINet-V2 [?] and PFNet  [?] employed a coarse-to-fine strategy.
+Inspired by natural predatory behavior, SINet-V2 [7] and PFNet  [25] employed a coarse-to-fine strategy.
 They first generated a preliminary location map for camouflaged objects and then refined it for segmentation.
 To enhance performance, several studies integrated auxiliary task into a joint learning framework.
-MGL [?] combined classification or boundary detection task with COD.
-ZoomNet [?] employed a zooming in and out strategy to the original inputs and processed appearance features at three different scales.
+MGL [26] combined classification or boundary detection task with COD.
+Liu et al.
+[27] investigated the part-object relationship to discover camouflaged objects.
+ZoomNet [28] employed a zooming in and out strategy to the original inputs and processed appearance features at three different scales.
+Then Pang et al.
+extended ZoomNet to ZoomNeXt [29], which simultaneously addresses image and video camouflaged object detection.
 Jia et al.
-[?] proposed the SegMaR framework, an iterative refinement approach designed to locate, magnify, and detect camouflaged objects.
-HitNet [?] elevated low-resolution representations by leveraging high-resolution features in an iterative feedback loop, effectively mitigating challenges such as edge blurring and detail degradation.
-Video-based COD.
+[30] proposed the SegMaR framework, an iterative refinement approach designed to locate, magnify, and detect camouflaged objects.
+Ji et al.
+[9] designed a two-branch framework to encode the context and texture of camouflage objects under gradient supervision.
+Huang et al.
+[31] designed progressively neighboring token enhanced decoder to exploit imperceptible cues for detecting camouflaged objects.
+Zhang et al.
+[32] introduced predictive uncertainty estimation framework to address model and data uncertainty simultaneously in camouflaged scenes .
+HitNet [33] elevated low-resolution representations by leveraging high-resolution features in an iterative feedback loop, effectively mitigating challenges such as edge blurring and detail degradation.
+Methods [34, 35] designed frameworks to mine the subtle cues of camouflaged objects in the frequency domain.
+Yao et al.
+[36] proposed a hierarchical graph interaction network to refine ambiguous regions.
+Hao et al.
+[37] proposed a simple yet effective general architecture for both COD and SOD tasks.
+II-B Video-based COD
 For the VCOD task, motion cues are crucial to camouflaged object detection.
 Bidau et al.
-[?] introduced a method by approximating various motion models derived from dense optical flow.
-Lamdouar et al.[?] introduced a video registration and segmentation network to detect camouflaged objects, employing optical flow and a difference image as inputs.
+[38] introduced a method by approximating various motion models derived from dense optical flow.
+Zhang et al.
+[39] proposed a camouflage modeling strategy and fused it with discriminative modeling in a Bayesian framework for moving object detection.
+Lamdouar et al.[11] introduced a video registration and segmentation network to detect camouflaged objects, employing optical flow and a difference image as inputs.
 However, the utilization of inaccurate optical flow may result in accumulative errors in mask prediction.
 To address this challenge, Cheng et al.
-[?] proposed a two-stage model that implicitly models and leverages motion information.
+[10] proposed a two-stage model that implicitly models and leverages motion information.
+Subsequently, to eliminate inaccuracies stemming from implicit motion modeling in SLT-Net [10], Hui et al.
+[40] introduced a motion-induced consistency preserving approach between frames with a feature pyramid framework.
+Lu et al.
+[41] introduced a weakly-supervised framework for VCOD.
+Hui et al.
+[42] leveraged temporal and spatial relationships between frames to generate prompts for SAM.
 Different from the above approaches, we propose to handle motion cues in an explicit way as aforementioned, and a two-stream architecture is designed to conduct both optical flow estimation and segmentation.
-Visual Prompt Learning.
-Recently, prompt learning has emerged as a new paradigm that has significantly enhanced the performance of natural language processing (NLP) tasks [?].
-Besides, the prompting paradigm has been adopted in many computer vision tasks [??].
-Work [?] modified transformer layers by introducing memory tokens, constituting a set of learnable embedding vectors.
-VPT [?] employed a similar strategy by applying learnable embedding vectors to transformer encoders, achieving noteworthy performance across various downstream recognition tasks.
-Based on this idea, ViPT [?] integrated modality-complementary visual prompts for task-oriented multi-modal tracking.
-The popular Segment Anything Model (SAM) [?] integrates various visual prompts like points, boxes, or masks to achieve tailored object segmentation and exhibits decent zero-shot generalization.
+II-C Video Object Segmentation (VOS) & Video Salient Object Detection (VSOD)
+VCOD can be seen as a specialized task of video object segmentation (VOS) that segments objects across consecutive video frames.
+Mei et al.
+[43] proposed a transformer-based framework to leverage temporal and spatial relationships across frames.
+To increase the detection speed, Park et al.
+[44] introduced a novel network capable of dynamically selecting mask generation methods, either by reusing features from prior frames or processing the entire network.
+To incorporate these advantages, memory-based networks [45, 46] have also been explored to better utilize historical information.
+These VOS networks use the current frame to query a memory bank storing historical features and corresponding object masks from past frames.
+On the other hand, the objective of video salient object detection (VSOD) stands in direct contrast to VCOD.
+VSOD focuses on locating and segmenting the most prominent objects from sequences.
+Chen et al.
+[47] utilized spatial and temporal cues along with local constraints to achieve global saliency optimization.
+Li et al.
+[48] proposed a motion-based bilateral network for background estimation, and the background estimation results are then merged with instance embeddings into a graph, where edges connect pixels across different frames for multi-frame reasoning.
+Song et al.
+[49] proposed a novel recurrent network to extract multi-scale spatial features, which are then concatenated and fed into an extended deep bidirectional ConvLSTM to learn spatiotemporal information.
+Cong et al.
+[50] designed a two-stage method: first, obtaining the spatial saliency of each frame through sparsity-based reconstruction, and then capturing the sequential correspondence in the temporal space via progressive sparsity-based propagation.
+Xu et al.
+[51] proposed a novel method for modeling motion energy based on four aspects: gradient flow field, motion direction, motion magnitude, and the spatial gradients of video frames.
+Yan et al.
+[52]  et al.
+leveraged optical flow estimation to generate pseudo-labels for some unannotated frames in the dataset, and further enhances the spatio-temporal correlation between video frames using Non-local [53] and ConvGRU [54].
+Chen et al.
+[55] integrated a lightweight temporal model into the spatial branch, coarsely locating spatial saliency regions associated with highly confident salient motion.
+Simultaneously, the spatial branch itself can iteratively optimize the temporal model in a multi-scale manner.
+Ji et al.
+[56] proposed a full-duplex strategy to obtain more stable consistent features.
+Zhao et al.
+[57] introduced a space-time memory-based network and leveraged high-level features to refine low-level details.
+Compared to existing VOS and VSOD methods, our model excels in discerning subtle differences between camouflaged objects and their highly similar surroundings by simultaneously integrating camouflage properties with explicit motion modeling information.
+II-D Visual Prompt Learning
+Recently, prompt learning has emerged as a new paradigm that has significantly enhanced the performance of natural language processing (NLP) tasks [58].
+Besides, the prompting paradigm has been adopted in many computer vision tasks [17, 18].
+Work [17] modified transformer layers by introducing memory tokens, constituting a set of learnable embedding vectors.
+VPT [18] employed a similar strategy by applying learnable embedding vectors to transformer encoders, achieving noteworthy performance across various downstream recognition tasks.
+Based on this idea, ViPT [19] integrated modality-complementary visual prompts for task-oriented multi-modal tracking.
+The popular Segment Anything Model (SAM) [59] integrated various visual prompts like points, boxes, or masks to achieve tailored object segmentation and exhibited decent zero-shot generalization.
 Previous researches have mainly focused on specific tasks like classification, tracking, or segmentation.
-In this paper, inspired by [?], we treat intermediate features from one stream as a prompt for injecting complementary information to the other, and propose an interactive prompting scheme for the VCOD task.
-3 Methodology
+In this paper, inspired by [19], we treat intermediate features from one stream as a prompt for injecting complementary information to the other, and propose an interactive prompting scheme for the VCOD task.
+III Methodology
 In this work, we propose an end-to-end trainable network EMIP to jointly optimize camouflaged object detection and optical flow estimation.
@@
 2.
-Note that, for object segmentation stream, a set of features $\{f_{n}^{i}\in\mathbb{R}^{H/{2^{i+1}}\times W/{2^{i+1}}\times C_{i}},n\in\{t,t-1\},i=1,...,4\}$ with different scales are extracted from a vision transformer backbone (i.e., PVTv2-B5 [?]), the same as SLT-Net [?], for fair comparison.
+Note that, for object segmentation stream, a set of features $\{f_{n}^{i}\in\mathbb{R}^{H/{2^{i+1}}\times W/{2^{i+1}}\times C_{i}},n\in\{t,t-1\},i=1,...,4\}$ with different scales are extracted from a vision transformer backbone (i.e., PVTv2-B5 [60]), the same as SLT-Net [10], for fair comparison.
 $W$, $H$, and $C$ represent the width, height, and channel number, respectively.
-Similar to [?], we adopt the top-three features ($f_{t}^{2},f_{t}^{3},f_{t}^{4}$) for appearance modeling, and discard the first-layer feature $f_{t}^{1}$.
-3.1 Fundamental Model for Motion Modeling
-To achieve more effective integration of motion information for the VCOD task, we select GMFlow [?] as our fundamental model, which is trained on $\sim$50k frame pairs for optical flow estimation.
+Similar to [61], we adopt the top-three features ($f_{t}^{2},f_{t}^{3},f_{t}^{4}$) for appearance modeling, and discard the first-layer feature $f_{t}^{1}$.
+Figure 2: Overall architecture of the proposed EMIP, which consists of two separate streams: explicit motion modeling stream (upper) and object segmentation stream (lower).
+We use GMFlow [16] as the fundamental model to handle motion cues.
+With the camouflage feeder and motion collector, segmentation and motion prompts are injected into each task-specific stream to compensate essential information.
+The fire/snowflake symbols indicate that the model parameters in this part or block are designated as learnable/frozen.
+III-A Fundamental Model for Motion Modeling
+To achieve more effective integration of motion information for the VCOD task, we select GMFlow [16] as our fundamental model, which is trained on $\sim$50k frame pairs for optical flow estimation.
 As shown in Fig.
@@
 It initially employs the CNN encoder to capture low-level features such as edges, colors, and textures.
-Subsequently, the transformer decoder, composed of a sequence of self- and cross-attention layers, predicts optical flow map $\boldsymbol{V}\in\mathbb{R}^{H\times W\times 2}$ and matching distribution $\boldsymbol{M}\in\mathbb{R}^{H\times W\times H\times W}$ simultaneously.
+Subsequently, the transformer decoder, composed of a sequence of self- and cross-attention layers, predicts optical flow map $\boldsymbol{V}\in\mathbb{R}^{H/8\times W/8\times 2}$ and matching distribution $\boldsymbol{M}\in\mathbb{R}^{H/8\times W/8\times H/8\times W/8}$ simultaneously.
 $\boldsymbol{V}$ is derived from $\boldsymbol{M}$ through operations such as pixel-grid sampling.
-Here, we flatten the first two dimensions of $\boldsymbol{M}$ to obtain $\boldsymbol{G}\in\mathbb{R}^{H\times W\times HW}$, and employ $\boldsymbol{G}$ as the motion prompt for interacting with segmentation features, leveraging its detailed pixel-to-pixel matching information.
-Additionally, the flow map $\boldsymbol{V}$ is utilized to reconstruct frame $I_{t}$ from $I_{t-1}$, leading to the computation of a self-supervised loss.
-For more details about our fundamental motion estimation model, we refer readers to GMFlow [?].
-3.2 Segmentation-to-Motion Prompt
+Here, we flatten the first two dimensions of $\boldsymbol{M}$ to obtain $\boldsymbol{G}\in\mathbb{R}^{H/8\times W/8\times HW/64}$, and employ $\boldsymbol{G}$ as the motion prompt for interacting with segmentation features, leveraging its detailed pixel-to-pixel matching information.
+Additionally, the optical flow map $\boldsymbol{V}$ is utilized to reconstruct frame $I_{t}$ from $I_{t-1}$, leading to the computation of a self-supervised loss.
+For more details about our fundamental motion estimation model, we refer readers to GMFlow [16].
+III-B Segmentation-to-Motion Prompt
 A good prompt can fully exploit the potential of the fundamental model, the same for our EMIP.
 Therefore, the segmentation prompt has to determine where and how to embed it in the fundamental model, considering the nature of the optical flow estimation task.
-Position of prompt.
+III-B1 Position of prompt
 The selection of the prompt position is guided by two key observations.
@@
 Thus, preserving the integrity of this architecture becomes paramount for the coherent generation of optical flow.
-Secondly, the architectural arrangement of neural networks involves a stratification where lower layers are dedicated to extracting basic features such as edges, colors, and textures, each exhibiting distinctive characteristics across various modalities [?].
-This stratification has been shown by [?] to be particularly advantageous for task like optical flow estimation, emphasizing the importance of low-level visual correspondences.
+Secondly, the architectural arrangement of neural networks involves a stratification where lower layers are dedicated to extracting basic features such as edges, colors, and textures, each exhibiting distinctive characteristics across various modalities [62].
+This stratification has been shown by [15] to be particularly advantageous for task like optical flow estimation, emphasizing the importance of low-level visual correspondences.
 Leveraging this insight, we position the segmentation prompt at the lower layers of both the segmentation and motion modeling streams, with the direction being from segmentation to motion.
-Further details are discussed in Section 4.3, where we delve into various prompt position choices and their impact on results.
-Camouflage feeder.
+Further details are discussed in Section IV-D, where we delve into various prompt position choices and their impact on results.
+III-B2 Camouflage feeder
 To better incorporate the segmentation prompt into the motion stream, we design camouflage feeder.
 As shown in Fig.
-2, our camouflage feeder is based on cross attention with a residual connection, which takes the flow input feature $f^{flow}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of query, and the segmentation prompt feature $f^{2}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of key and value.
+2 right, our camouflage feeder is based on cross attention with a residual connection, which takes the optical flow input feature $f^{flow}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of query, and the segmentation prompt feature $f^{2}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of key and value.
 The residual connection is to maintain more query-related cues, and $d$ represents the embedding dimension.
@@
 and $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ represent query, key, and value matrices, respectively.
-$\bf Q/\bf K$ and $\bf V$ are derived by applying layer normalization and 3$\times$3 convolution to the input feature maps $f^{flow}_{t}$ and $f^{2}_{t}$, respectively.
-Here, we employ a Gated-Dconv Feed-forward Network (GDFN) [?] to suppress the aggregation of noise prompt.
-The GDFN process is defined as:
+$\bf Q/\bf K$ and $\bf V$ are derived by applying layer normalization and 3$\times$3 convolution to the input $f^{flow}_{t}$ and $f^{2}_{t}$, respectively.
+Here, we employ a Gated-Dconv Feed-forward Network (GDFN) [63] to suppress the aggregation of noisy prompt, and GDFN for input $\mathbf{X}$ is defined as:
 $\mathtt{GDFN}(\mathbf{X})=\mathtt{C1}\left[\phi(\mathtt{C3}(\mathtt{C1}(\mathtt{LN}(\mathbf{X}))))\odot\mathtt{C3}(\mathtt{C1}(\mathtt{LN}(\mathbf{X})))\right],$ (4)
 where $\mathtt{C1}$ and $\mathtt{C3}$ are 1$\times$1 convolution and 3$\times$3 depth-wise convolution, respectively.
-$\phi$ is GELU function [?] and $\odot$ denotes element-wise multiplication.
+$\phi$ is GELU function [64] and $\odot$ denotes element-wise multiplication.
 Note that, $\mathtt{GDFN}$ comprises two separate paths, one of which is activated with GELU function.
 Subsequently, an element-wise product is applied between these two paths.
-Literature [?] has demonstrated that the gating mechanism in GDFN can control information flow and yield better performance compared with the conventional feed-forward network (FN) [?] for feature restoration.
-For the same purpose, to mitigate the impact of the integrated noise prompts, we employ GDFN to control noise information flow.
-Finally, the resulting segmentation-prompted motion features $\hat{f}^{flow}_{t}$ with both knowledges of motion and appearance can be well adapted to subsequent motion reconstruction.
-Prompt learning.
-To better learn the motion information, inspired by [?], we employ a self-supervised loss to optimize the motion-to-segmentation prompt learning.
-Further elaboration is provided in Section 3.4.
-3.3 Motion-to-Segmentation Prompt
+Literature [63] has demonstrated that the gating mechanism in GDFN can control information flow and yield better performance compared with the conventional feed-forward network (FN) [65] for feature restoration.
+For the same purpose, to mitigate the impact of the integrated noisy prompts, we employ GDFN to control noisy information flow.
+Finally, the resulting segmentation-prompted motion features $\hat{f}^{flow}_{t}$ with both knowledge of motion and appearance can be well adapted to subsequent motion reconstruction.
+III-B3 Prompt learning
+To better learn the motion information, inspired by [66], we employ a self-supervised loss to optimize the motion-to-segmentation prompt learning. Further elaboration is provided in Section III-D.
+III-C Motion-to-Segmentation Prompt
 Motion-to-segmentation prompt formally provides an auxiliary motion flow that is both temporally and spatially coordinated with the segmentation stream.
@@
 Thus, we introduce another module named Motion Collector to integrate motion prompts into the segmentation stream.
-Given its functional equivalence of motion collector with the camouflage feeder, we opt to maintain the structural configuration of the interaction block rather than introducing a redesign.
-Specifically, upon receiving motion feature $\boldsymbol{G}$ (flattening the first two dimensions of $\boldsymbol{M}$), we commence by applying a 3$\times$3 convolution operation to $\boldsymbol{G}$ to adjust the channel dimensions to align with $f^{2}_{t}$.
+Given the functional analogy of motion collector with the camouflage feeder, we opt to maintain the structural configuration of the interaction block rather than introducing a redesign (as illustrated in Fig.
+2 right).
+Specifically, upon receiving motion feature $\boldsymbol{G}$ (flattening the first two dimensions of $\boldsymbol{M}$), we then apply a 3$\times$3 convolution operation to $\boldsymbol{G}$ to adjust and reduce its channel dimension to align with $f^{2}_{t}$.
 Then we feed the motion-to-segmentation prompt $\hat{\boldsymbol{G}}=\mathtt{Conv}(\boldsymbol{G})$ and $f^{2}_{t}$ to the motion collector to obtain the prompted appearance feature $f_{t\leftarrow t-1}^{2}$.
-Subsequently, following that most VCOD works [??] use neighbor connection decoder (NCD) [?] to predict the segmentation map, we use NCD to decode the appearance features.
-NCD with camouflage feeder can be regarded as the prediction head for the motion fundamental model, similar to the concepts of some previous prompt learning methods [??] that design an extra prediction head after freezing the backbone parameters.
+Subsequently, following that most VCOD works [9, 10] use neighbor connection decoder (NCD) [10] to predict the segmentation map, we use NCD to decode the appearance features.
+NCD with motion collector can be regarded as the prediction head for the motion fundamental model, similar to the concepts of some previous prompt learning methods [9, 10] that design an extra prediction head after freezing the backbone parameters.
 Specifically, the appearance feature $f_{t\leftarrow t-1}^{2}$ prompted by motion combined with $f^{3}_{t}$ and $f^{4}_{t}$ are fed to NCD to obtain the predicted map.
 The optimization of motion-to-segmentation prompt and prediction head is under the constraint of segmentation loss, and more details will be given in Sec.
-3.4.
+III-D.
 Discussion.
-The motivation for our prompt learning strategy, freezing the motion stream and fully tuning the segmentation stream, is from the observed performance degradation of the PVT [?] pre-trained on ImageNet when it is directly applied to the VCOD task.
-Similar challenges are also encountered with large models like SAM [?], which struggle in camouflage scenes [?].
+The motivation for our prompt learning strategy, freezing the motion stream and fully tuning the segmentation stream, is from the observed performance degradation of the PVT [60] pre-trained on ImageNet when directly applied to the VCOD task.
+Similar challenges are also encountered with large models like SAM [59], which struggle in camouflaged scenes [67].
 Currently, there lacks a freezable foundational model that excels on the VCOD task.
 In contrast, we find that the optical flow model could achieve good generalization in most VCOD scenarios, when most of its parameters are frozen.
-3.4 Supervision and Loss Function
+III-D Supervision and Loss Function
 To ensure optimization for each component of the model, we separately define loss functions for the motion and segmentation stream, and then use these two losses to optimize the entire model jointly.
 Considering the lack of ground truth (GT) optical flow in video camouflaged scenarios, we propose a self-supervised strategy to learn the optical flow estimation stream.
-For the sake of brevity and clarity, we refer to [?] and warp the frame $I_{t-1}$ according to the optical flow estimation $\boldsymbol{V}$ to obtain the reconstructed reference frame for $I_{t}$, which is denoted as $\hat{I}_{t}$, i.e., $\hat{I}_{t}=\mathtt{Warp}(\boldsymbol{V},I_{t-1})$.
+For the sake of brevity and clarity, we refer to [66] and warp the frame $I_{t-1}$ according to the optical flow estimation $\boldsymbol{V}$ to obtain the reconstructed reference frame for $I_{t}$, which is denoted as $\hat{I}_{t}$, i.e., $\hat{I}_{t}=\mathtt{Warp}(\boldsymbol{V},I_{t-1})$.
 We supervise the optical flow by computing the distance between $I_{t}$ and $\hat{I}_{t}$.
@@
 $\mathcal{L}_{\text{flow}}=\mathtt{SSIM}(I_{t},\hat{I}_{t}),$ (5)
-where $\mathtt{SSIM}$ is pixel-wise photometric loss [?], which is commonly used for reflecting image distortion from three aspects: brightness, contrast, and structure.
-For the segmentation stream, we employ a hybrid loss function [?], the same as SLT-Net [?].
-The hybrid loss $\mathcal{L}_{\text{seg}}$ can be defined as $\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{IoU}}+\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{e-loss}}$, which includes IoU loss, binary cross-entropy loss, and enhanced-alignment loss.
-Finally, we jointly optimize the motion modeling and object segmentation streams, whose total loss is formulated as:
+where $\mathtt{SSIM}$ is pixel-wise photometric loss [68], which is commonly used for reflecting image distortion from three aspects: brightness, contrast, and structure.
+For the segmentation stream, we employ a hybrid loss function [69], the same as SLT-Net [10].
+The hybrid loss $\mathcal{L}_{\text{seg}}$ is defined as $\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{IoU}}+\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{e-loss}}$, which includes IoU loss, binary cross-entropy loss, and enhanced-alignment loss.
+Finally, we jointly optimize the motion modeling and object segmentation streams, and the total loss is formulated as:
 $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{seg}}+\mathcal{L}_{\text{flow}}.$ (6)
-3.5 Long-term Consistency Modeling
-Figure 3: Long-term modeling scheme (${}^{\dagger}$).
+III-E Long-term Consistency Modeling
+Figure 3: Overview of our long-term modeling scheme (EMIP). EMIP consists of a frozen EMIP and other five learning modules (i.e., Memory Encoder, Query Encoder, STM, Motion Collector, and NCD).
 The short-term model captures correlated information between adjacent frames, but due to the impact of noises and the limitation of motion estimation model, the short-term model may be less robust.
@@
 Fig.
-3 illustrates our long-term consistency modeling scheme, termed ${}^{\dagger}$.
-Specifically, during training, we freeze the short-term model after it converges and then add the long-term memory module for further learning.
+3 illustrates our long-term consistency modeling scheme, termed EMIP.
+Specifically, during training, we freeze the short-term model (referring to the explicit motion modeling stream and the object segmentation stream in Fig.
+3) after it converges and then add the long-term memory module for further learning.
 Given a video sequence {${{I}_{1},...,{I}_{t-1},{I}_{t}},t>1$}, we sequentially input current frame ${I}_{t}$ and preceding frame ${I}_{t-1}$ into the short-term model to obtain the appearance features $f^{2}_{t}$ and the motion-to-segmentation prompt $\hat{\boldsymbol{G}}$ between the two frames.
-The element-wise sum of $f^{2}_{t}$ and $\hat{\boldsymbol{G}}$ is conducted, and the results are then fed to an encoder consisting of two convolutional layers.
-The processed information is later stored into the memory pool.
+The element-wise sum of $f^{2}_{t}$ and $\hat{\boldsymbol{G}}$ is conducted, and the results are then fed to the memory encoder to be mapped as key $K_{Mt}$ and value $V_{Mt}$.
+The memory encoder consists of a convolutional layer, followed by a normalization layer and a ReLU activation function, and then two parallel convolutional layers.
+The input features are processed to obtain the key-value mappings $K_{Mt}$ and $V_{Mt}$.
+The processed key-value mappings are later stored into the memory pool.
 Considering the computational burden and minor impact of distant frames, we implement a FIFO (First-In First-Out) memory pool with fixed capacity ${L}$.
 Note that such a fixed capacity is configurable according to practical applications.
-Referring to the long-term setting of SLT-Net, we set ${L}$ to 5 in this paper.
-Since the memory pool stores appearance and motion features for both the current and historical frames, the features extracted from the current frame are then employed to query the memory pool using a space-time memory read block (STM) [?], yielding the long-term prompt ${P}_{L}$.
-Next, ${P}_{L}$ and $f^{2}_{t}$ are fed to the motion collector (re-use the same motion collector of the short-term period) for interaction, and the output of collector is subsequently fed to NCD for decoding.
+Referring to the long-term setting of SLT-Net [10], we set ${L}$ to 5 in this paper.
+The object feature $f^{2}_{t}$ extracted from the current frame are fed to the query encoder, which consists of two parallel convolutional layers used to compute the query key-value mappings $K_{Q}$ and $V_{Q}$.
+These mappings are then used to query the memory pool, which stores appearance and motion features for both the current and historical frames, via a space-time memory read block (STM) [46] (referring to the right part of Fig.
+3).
+This process yields the long-term prompt ${P}_{L}$.
+Specifically, the key-value mappings of each historical frame in the memory pool are concatenated along the temporal dimension to form $K_{M}$ and $V_{M}$.
+These concatenated mappings are then interacted with the query key-value mappings $K_{Q}$ and $V_{Q}$ through the STM module, ultimately producing the query results.
+Next we apply a $3\times 3$ convolution operation to ${P}_{L}$ to adjust and reduce its channel dimension to align with $f^{2}_{t}$.
+Then, $\hat{P}_{L}=\mathtt{Conv}({P}_{L})$ and $f^{2}_{t}$ are fed to the motion collector [^1] Can re-use the same motion collector of the short-term period if the short-term output is no longer needed.
+for interaction, and the output of collector, denoted as $f_{t\leftarrow t-L}$, is subsequently fed to NCD for decoding.
 We still adopt the hybrid loss $\mathcal{L}_{\text{seg}}$, the same as the short-term version, for training the long-term counterpart.
@@
 Hence, our scheme is theoretically more suitable for real-time and practical applications, where only past information is available.
-4 Experiments
-4.1 Implementation Details and Datasets
-For fair comparisons, following the training strategy with [?], we train our model with a two-stage pipeline: first train the backbone on the static training set of COD10K (3,040 images) [?], and then fine-tune the whole model with the temporal components on the training set of MoCA-Mask (19,313 frames) [?].
-Evaluation is performed on two widely recognized VCOD benchmarks: CAD (836 frames) [?] and MoCA-Mask test set (3,626 frames).
-Five evaluation metrics are adopted to assess model performance, i.e., structure measure ($\mathcal{S}_{\alpha}$[?], weighted F-measure ($F_{\beta}^{w}$[?], mean absolute error ($\mathcal{M}$), and mean value of Dice and IoU.
-The proposed EMIP is implemented by PyTorch [?] on a single NVIDIA 4090 GPU.
-During the training stage, the batch size is set to 6, and the maximum training epoch is 60.
-The entire model is optimized using the Adam optimizer [?] with a cosine annealing strategy.
+IV Experiments and Results
+IV-A Datasets and Metrics
+IV-A1 Datasets
+Following [10, 40], we conduct experiments on two widely recognized VCOD benchmarks: MoCA-Mask [10] and CAD [38].
+Among them, MoCA-Mask stands as the most challenging dataset, comprising 19,313 frames across 71 clips for training, and 3,626 frames of 16 clips for testing.
+The CAD dataset includes 836 frames of 9 clips designated for testing.
+To assess the generalizability of our model, we conducted experiments on four widely used VSOD/VOS datasets: DAVIS16 [70] (with 30 training clips and 20 testing clips), FBMS [71] (29 training clips and 30 testing clips), ViSal [72] (comprising 17 video sequences for testing) and SegV2 [73] (including 13 clips for testing).
+IV-A2 Evaluation Metrics
+We adopt widely recognized evaluation metrics to assess our model performance, namely: structure measure ($\mathcal{S}_{\alpha}$[74], weighted F-measure ($F_{\beta}^{w}$[75], F-measure (max $\mathcal{F}$[76], mean absolute error ($\mathcal{M}$[77], and mean value of Dice and IoU.
+These metrics provide a comprehensive and reliable assessment of model performance.
+IV-B Implementation Details
+For fair comparisons, we adhere to the training settings outlined in [10] and employ PVTv2 [60] as the feature extraction backbone.
+All input images are resized to 352$\times$352 and subject to data augmentation techniques, including color enhancement and random rotation.
+A two-stage training pipeline in [10] is also adopted by our model: first train the backbone on the static training set of COD10K (3,040 images) [7], and then fine-tune the whole model with the temporal components on the training set of MoCA-Mask (19,313 frames) [10].
+The entire model is optimized using the Adam optimizer [78] with a cosine annealing strategy.
 The maximum and minimum learning rates, along with the maximum adjusted iterations, are set to 1e-5, 1e-6, and 20, respectively.
-All the input images are resized to 352$\times$352 and subjected to data augmentation techniques, including color enhancement and random rotation.
+EMIP is trained for 7 hours over 60 epochs on an NVIDIA 4090 GPU with 16,536 MB of memory using a batch size of 6.
 For inference, EMIP takes any two consecutive frames as input and outputs the segmentation prediction together with the corresponding optical flow estimation.
-However, for ${}^{\dagger}$, it requires consecutive-frame pairs to be input sequentially, in order to leverage its long-term property.
+However, for EMIP, it requires consecutive-frame pairs to be input sequentially, in order to leverage its long-term property.
+For VSOD, we train our model on the training set of DAVIS16 (30 clips) and FBMS (29 clips) as [56].
+The proposed EMIP is implemented by PyTorch [79], and all experiments are conducted on an NVIDIA RTX 4090 GPU.
 […]
@@
 […]
-Image-based SINet [?] CVPR’20 0.598 0.231 0.028 0.276 0.202 0.636 0.346 0.041 0.381 0.283
-SINet-V2 [?] TPAMI’22 0.588 0.204 0.031 0.245 0.180 0.653 0.382 0.039 0.413 0.318
-DGNet [?] MIR’23 0.581 0.184 0.024 0.222 0.156 0.686 0.416 0.037 0.456 0.340
-FEDER [?] CVPR’23 0.560 0.165 0.031 0.194 0.137 0.691 0.444 0.029 0.474 0.375
-HitNet [?] AAAI’23 0.623 0.299 0.019 0.318 0.254 0.685 0.463 0.031 0.478 0.373
-Video-based RCRNet [?] ICCV’19 0.555 0.138 0.033 0.171 0.116 0.627 0.287 0.048 0.309 0.229
-MG [?] ICCV’21 0.530 0.168 0.067 0.181 0.127 0.594 0.336 0.059 0.368 0.268
-PNS-Net [?] MICCAI’21 0.544 0.097 0.033 0.121 0.101 0.655 0.325 0.048 0.384 0.290
-SLT-Net [?] CVPR’22 0.637 0.304 0.027 0.356 0.271 0.696 0.471 0.031 0.484 0.392
-${}^{\dagger}$ [?] CVPR’22 0.631 0.311 0.027 0.360 0.272 0.696 0.481 0.030 0.493 0.401
-EMIP (Ours) 0.669 0.374 0.017 0.424 0.326 0.710 0.504 0.029 0.528 0.415
-${}^{\dagger}$ (Ours) 0.675 0.381 0.015 0.426 0.333 0.719 0.514 0.028 0.536 0.425
-Figure 4: Visual comparisons of our models (EMIP and ${}^{\dagger}$) with eight state-of-the-art methods.
-4.2 Comparisons with State-of-the-arts
-To demonstrate the effectiveness of our method, we compare our model with 10 cutting-edge methods, comprising five image-based and five video-based.
-To ensure fair evaluation, predictions from these methods are either directly downloaded from the official website or generated using models retrained with open-source code.
-Quantitative results are presented in Table 1.
-Notable performance gains can be observed of EMIP and ${}^{\dagger}$ over recently proposed image-based COD and VCOD models, like HitNet [?] and SLT-Net [?], across all five evaluation metrics.
-Several visual samples are shown in Fig.
-4.
-EMIP appears to be more effective at simultaneously conducting motion and segmentation, and incorporating both enhanced features further makes it better for detecting camouflaged objects.
-Additionally, predictions of consecutive frames are listed in Fig.
+Image-based EGNet [80] ICCV’19 VGG-16 0.547 0.110 0.035 0.143 0.096 0.619 0.298 0.044 0.324 0.243
+BASNet [81] CVPR’19 ResNet-34 0.561 0.154 0.042 0.190 0.137 0.639 0.349 0.054 0.393 0.293
+CPD [82] CVPR’19 VGG-16 0.561 0.121 0.041 0.162 0.113 0.622 0.289 0.049 0.330 0.239
+PraNet [61] MICCAI’20 Res2Net-50 0.614 0.266 0.030 0.311 0.234 0.629 0.352 0.042 0.378 0.290
+SINet [8] CVPR’20 ResNet-50 0.598 0.231 0.028 0.276 0.202 0.636 0.346 0.041 0.381 0.283
+SINet-V2 [7] TPAMI’22 ResNet-50 0.588 0.204 0.031 0.245 0.180 0.653 0.382 0.039 0.413 0.318
+ZoomNet [28] CVPR’22 ResNet-50 0.582 0.211 0.033 0.224 0.167 0.587 0.225 0.063 0.246 0.166
+DGNet [9] MIR’23 PVT 0.581 0.184 0.024 0.222 0.156 0.686 0.416 0.037 0.456 0.340
+FEDER [34] CVPR’23 ResNet-50 0.560 0.165 0.031 0.194 0.137 0.691 0.444 0.029 0.474 0.375
+FSPNet [31] CVPR’23 ViT 0.594 0.182 0.044 0.238 0.167 0.539 0.220 0.145 0.309 0.212
+PUENet [32] TIP’23 ViT 0.594 0.204 0.037 0.302 0.212 0.673 0.427 0.034 0.499 0.389
+HitNet [33] AAAI’23 PVT 0.623 0.299 0.019 0.318 0.254 0.685 0.463 0.031 0.478 0.373
+FSEL [83] ECCV’24 PVT 0.596 0.260 0.053 0.219 0.151 0.649 0.368 0.053 0.434 0.325
+HGINet [36] TIP’24 ViT 0.610 0.251 0.030 0.303 0.221 0.680 0.437 0.050 0.501 0.392
+Video-based RCRNet [52] ICCV’19 ResNet-50 0.555 0.138 0.033 0.171 0.116 0.627 0.287 0.048 0.309 0.229
+MG [13] ICCV’21 VGG-style 0.530 0.168 0.067 0.181 0.127 0.594 0.336 0.059 0.368 0.268
+PNS-Net [84] MICCAI’21 Res2Net-50 0.544 0.097 0.033 0.121 0.101 0.655 0.325 0.048 0.384 0.290
+SLT-Net [10] CVPR’22 PVT 0.637 0.304 0.027 0.356 0.271 0.696 0.471 0.031 0.484 0.392
+SLT-Net [10] CVPR’22 PVT 0.631 0.311 0.027 0.360 0.272 0.696 0.481 0.030 0.493 0.401
+IMEX [40] TMM’24 ResNet-50 0.661 0.371 0.020 0.409 0.319 0.684 0.452 0.033 0.469 0.370
+EMIP (Ours) PVT 0.669 0.374 0.017 0.424 0.326 0.710 0.504 0.029 0.528 0.415
+EMIP (Ours) PVT 0.675 0.381 0.015 0.426 0.333 0.719 0.514 0.028 0.536 0.425
+Figure 4: Visual comparisons of our models (EMIP and EMIP) with eight state-of-the-art methods.
+We select some difficult scenarios, including dusky night, fast-moving objects, stationary objects, small objects, and noisy backgrounds.
+IV-C Comparisons with State-of-the-arts
+To demonstrate the effectiveness of our EMIP, we compare it with various state-of-the-arts.
+These compared methods can be categorized into two types: (i) Image-based camouflaged object detection methods [80, 81, 82, 61, 8, 7, 28, 9, 34, 31, 32, 33], which are designed for detecting objects in static camouflaged scenes; (ii) Video-based object detection methods [52, 13, 84, 10, 40], which focus on identifying camouflaged or moving objects in dynamic video sequences.
+Notably, we include comparisons with SLT-Net [10] and IMEX [40], two most relevant state-of-the-art approaches for video camouflaged object detection.
+For fair evaluation, the predictions from these methods are either directly downloaded from the official repositories or generated using open-source code provided by the original authors.
+IV-C1 Quantitative Evaluation
+Table I presents the experimental results on VCOD datasets.
+Notably, on MoCA-Mask dataset, our method demonstrates significant improvements: (i) It surpasses the previous state-of-the-art image-based approach, HitNet [33], by 27% in $F_{\beta}^{w}$, underscoring the efficacy of incorporating motion cues into the VCOD task.
+(ii) Additionally, when compared to SLT-Net [10], which utilizes the same backbone as ours, our method shows a marked improvement, achieving a 20% gain in $F_{\beta}^{w}$.
+Although our model was trained on the MoCA-Mask dataset, the testing results on the CAD dataset underscore its superior robustness and quantitative performance enhancements.
+The notable performance gains of EMIP over existing and recently proposed techniques emphasize that the combination of explicit motion modeling and prompt learning substantially enhances the completeness of detected camouflaged objects.
+Additionally, by leveraging long-term temporal information, our model achieves state-of-the-art performance across all five evaluation metrics on both VCOD benchmarks.
+This indicates that preserving long-term consistency effectively suppresses motion noise of camouflaged objects and enhances the stability of video predictions.
+Compared to SLT-Net [10] and IMEX [40], which also utilizes long-term consistency information as external cues to refine predictions, our long-term strategy yields notable performance gains.
+It is worth noting that, unlike SLT-Net, which frames long-term modeling as a sequence-to-sequence problem (requiring all frames of a clip), our approach utilizes only historical frames relative to the current frame.
+This design makes our scheme theoretically more suitable for real-time and practical applications, where only past information is accessible.
+IV-C2 Qualitative Comparison
+Fig.
+4 illustrates qualitative comparisons by visualizing the segmentation results of several examples.
+Our model demonstrates superior alignment with ground truth, showcasing its enhanced capability in identifying camouflaged objects compared to other methods.
+Furthermore, the segmentation results over consecutive frames of the same clip are shown in Fig.
+5.
+One can see that incorporating the long-term modeling scheme, namely EMIP, can reduce errors in short-term prediction and lead to boosted performance.
+Fig.
+11 presents visual comparisons of optical flow prediction in camouflaged scenarios.
+The designed prompting strategy in our model ensures more precise responses within camouflaged regions, while minimizing the influence of irrelevant motion noise.
+Consequently, our model excels in concurrently performing the motion modeling and segmentation tasks, leveraging the integration of enhanced features to improve the detection of camouflaged objects.
+Figure 5: Visual comparisons on consecutive video frames.
+IV-D Ablation Studies
+To evaluate the effectiveness of each core component, we conduct thorough ablation studies by removing or substituting components from the complete EMIP.
+IV-D1 Camouflage feeder and motion collector
+As shown in Table II, we first validate the role of the two prompt integration modules, i.e., camouflage feeder and motion collector.
+We start with the baseline model (#1), which directly uses the segmentation stream to extract single-frame information for prediction, i.e., removing the motion modeling stream of EMIP.
+Configuration #2 means using only motion information to prompt the segmentation stream, meanwhile discarding segmentation-to-motion prompt.
+Model #3 is the full EMIP.
+Comparing configuration #2 with #1, it is evident that camouflaged objects are challenging to detect without inter-frame motion information.
+In scenes with movement, motion information provides additional context that can improve the accuracy of segmentation.
+It helps in identifying and separating camouflaged objects that are in motion, which might be challenging using appearance-based methods alone.
+Comparing configuration #3 with #2, we observe that incorporating the segmentation-to-motion prompt via the camouflage feeder into the motion stream enhances the robustness of motion cues.
+This improvement in motion cue robustness subsequently leads to superior performance in camouflaged object detection.
+The visualization results are presented in Fig.
 6.
-One can see that incorporating the long-term modeling scheme, namely ${}^{\dagger}$, can further refine results and lead to boosted performance.
-Table 2: Ablation results on camouflage feeder (CF) and motion collector (MC) modules of EMIP on MoCA-Mask dataset. The best results are highlighted in bold.
+Relying solely on appearance features makes it challenging to accurately detect or localize truly camouflaged objects in certain scenarios.
+With the introduction of the motion collector, the camouflaged object is detected using motion information.
+Furthermore, incorporating the camouflage feeder to inject camouflage priors effectively reduces interference and enables more robust localization of the camouflaged object.
+TABLE II: Ablation results on camouflage feeder (CF) and motion collector (MC) modules of EMIP on MoCA-Mask dataset. The best results are highlighted in bold.
 […]
@@
 3 0.669 0.374 0.017 0.424 0.326
-4.3 Ablation Studies
-We conduct thorough ablation studies by removing or substituting components from the complete EMIP.
-Camouflage feeder and motion collector.
-We first validate the role of the two prompt integration modules, i.e., camouflage feeder and motion collector.
-As shown in Table 2, we start with the baseline model (#1), which directly uses the segmentation stream to extract single-frame information for prediction, i.e., removing the motion modeling stream of EMIP.
-Configuration #2 means using only motion information to prompt the segmentation stream, meanwhile discarding segmentation-to-motion prompt.
-Model #3 is the full EMIP.
-The integration of motion information is beneficial for the detection of camouflaged objects in video sequences.
-By incorporating the segmentation-to-motion prompt through the camouflage feeder into the motion stream, we enhance the robustness of motion cues, further resulting in improved performance for segmentation.
-Full-tuning v.s.
-Freezing.
+Figure 6: Qualitative results of CF and MC.
+The red box indicates the ground-truth location of the camouflaged object, while the yellow box represents detected noise.
+With the incorporation of MC, the camouflaged object can be effectively identified.
+Moreover, integrating CF further suppresses noise.
+IV-D2 Full-tuning v.s. Freezing
 Existing models usually address motion modeling by fully tuning all the parameters.
 In contrast, our EMIP adopts a prompt learning strategy to conduct motion modeling by freezing the motion fundamental model.
-From the comparisons in Table 3, where both settings are preloaded with pre-trained weights, the superiority of freezing the motion model over full-tuning is evident.
-This implies the prompt learning strategy for the motion modeling stream can better exploit its potential on limited VCOD data.
-Table 3: Different learning strategies for the motion stream.
+From the comparisons in Table III, where both settings are preloaded with pre-trained weights, the superiority of freezing the motion model over full-tuning is evident.
+As can be seen from Fig.
+7, full-tuning the motion stream disrupts its ability to effectively model motion, introducing additional noise into camouflaged object prediction.
+Freezing all motion modeling layers allows the model to leverage robust and general features learned from large-scale motion estimation datasets during pre-training.
+This strategy ensures that the prompting process concentrates on learning camouflage-specific features, thereby enhancing the efficiency and effectiveness of the optimization process of our EMIP.
+As demonstrated in Table III, the prompt learning strategy within our framework for the motion modeling stream significantly exploits its potential on limited VCOD data, even in the absence of ground truth for optical flow.
+TABLE III: Quantitative comparison of full-tuning and freezing the explicit motion modeling stream.
 […]
@@
 Freezing (Ours) 0.669 0.374 0.017 0.424 0.326
-Table 4: Different prompt destinations on the motion stream.
+Figure 7: Qualitative results of full-tuning and freezing the motion stream. The yellow boxes represent false positive predictions.
+IV-D3 Prompt destination on the motion stream
+The prompt destination is crucial as it determines where to introduce the appearance prompt within the motion fundamental model.
+In this experiment, we selected $f^{2}_{t}$ to serve dual roles: as the segmentation-to-motion prompt and as the input for the motion collector.
+Subsequently, we tested its effectiveness by prompting various convolutional neural network (CNN) layers within the model.
+The specific layers chosen for this evaluation ranged from initial to deeper CNN layers, allowing us to observe the impact at different stages.
+The evaluation results shown in Table IV reveal that prompting with the features immediately after CNN Layer3 yields the best performance among all tested positions.
+Fig.
+8 shows that injecting camouflage priors as prompts after CNN Layer3, but before motion relationship modeling, effectively mitigates the interference of inaccurate motion noise.
+This observation suggests that the features extracted by CNN Layer3 strike an optimal balance between low-level details and high-level abstractions.
+The superior performance can be attributed to the necessity for distinct-level features to be complemented by corresponding prompts that enhance level-specific representations.
+Essentially, the matching prompt at this stage helps in refining and emphasizing the critical features pertinent to motion and appearance, thereby boosting the overall performance of the model.
+By ensuring that the prompt is aligned with the level-specific characteristics of the features, our model can more effectively capture and utilize the nuanced information present at this stage.
+This experiment illustrates the importance of precise prompt positioning in the context of motion and appearance integration.
+TABLE IV: Quantitative comparison of different prompt destinations on the motion stream.
 […]
@@
 CNN Layer3 0.669 0.374 0.017 0.424 0.326
-Prompt destination on the motion stream.
-The prompt destination should address where to feed the appearance prompt on the motion fundamental model.
-In this experiment, we utilize $f^{2}_{t}$ as both segmentation-to-motion prompt and motion collector input.
-Then we use it to prompt different CNN layers, and the evaluation results are presented in Table 4.
-It can be seen that prompting with the features after CNN Layer3 achieves superior performance.
-The underlying reason is that distinct-level features need the corresponding matching prompt to enhance level-specific representations.
-Prompt source from the segmentation stream.
-To explore the most suitable appearance prompt to feed to the motion stream, we conduct another experiment with different prompt features, i.e., $f^{2}_{t}$, $f^{3}_{t}$, and $f^{4}_{t}$.
-The experimental results are presented in Table 5, where it is evident that utilizing appearance features $f^{2}_{t}$ as prompt yields the best performance.
-One possible explanation for this phenomenon is that slight motion of a moving object is likely to be missed in lower-resolution feature maps, namely $f^{3}_{t}$ and $f^{4}_{t}$, and therefore degrading the efficacy for prompting to the motion stream.
-Table 5: Different prompt sources from the segmentation stream.
+Figure 8: Visualization of different prompt destinations on the motion stream. The red box indicates the ground-truth location of the camouflaged object.
+IV-D4 Prompt source from the segmentation stream
+To thoroughly investigate the most suitable appearance prompt to feed to the motion stream, we conducted an additional set of experiments using various prompt features, i.e., $f^{2}_{t}$, $f^{3}_{t}$, and $f^{4}_{t}$.
+The experimental results are summarized in Table V and Fig.
+9 , clearly demonstrating that employing appearance features $f^{2}_{t}$ as the prompt leads to superior performance.
+A plausible explanation for this observation is that subtle movements of objects in motion are likely to be overlooked in lower-resolution feature maps, such as $f^{3}_{t}$ and $f^{4}_{t}$.
+Consequently, this omission diminishes the efficacy of these features when used as prompts in the motion stream, thereby underscoring the importance of higher-resolution appearance features in capturing fine-grained motion details.
+TABLE V: Different prompt sources from the segmentation stream.
 […]
@@
 $f^{4}_{t}$ 0.624 0.289 0.023 0.331 0.251
-Prompt destination on the segmentation stream.
-We leverage motion information to selectively prompt distinct appearance features or simultaneously prompt all features by introducing additional motion collectors.
-The results, as presented in Table 6, demonstrate that utilizing motion to prompt feature $f^{2}_{t}$ outperforms other features ($f^{3}_{t}$ or $f^{4}_{t}$) or all features collectively.
-This implies that semantic features at corresponding hierarchical levels between segmentation and motion tasks can mutually enhance the representation of the original features.
-Yet, the integration of features from disparate levels, which lack alignment, may compromise the fidelity of the original feature representation.
-Table 6: Different prompt destinations on the segmentation stream.
+Figure 9: Qualitative results of different prompt sources from the segmentation stream.
+The red box denotes the ground-truth location of the camouflaged object, while the yellow box highlights the detected noise.
+IV-D5 Prompt destination on the segmentation stream
+We harness motion information to selectively prompt distinct appearance features or to simultaneously prompt all features by incorporating additional motion collectors into the model.
+The empirical results, as presented in Table VI, reveal that utilizing motion to specifically prompt feature $f^{2}_{t}$ yields superior performance compared to prompting either $f^{3}_{t}$ or $f^{4}_{t}$ individually or all features collectively.
+Fig.
+10 illustrates that applying a prompt on $f^{2}_{t}$ enables the model to focus on the most relevant object features while effectively suppressing background noise.
+However, when prompting all features, including $f^{2}_{t}$, mismatched prompt locations can introduce interference with the original features, ultimately degrading detection performance.
+This finding suggests that semantic features at corresponding hierarchical levels between segmentation and motion tasks can synergistically enhance the representation of the original features.
+However, integrating features from different hierarchical levels, which lack proper alignment, may compromise the integrity and fidelity of the original feature representation, thereby negatively affecting overall performance.
+TABLE VI: Different prompt destinations on the segmentation stream.
 […]
@@
 0.639 0.313 0.020 0.361 0.271
-Effectiveness of motion self-supervision.
+Figure 10: Qualitative results of different prompt destinations on the segmentation stream. The yellow boxes represent false positive detections.
+IV-D6 Effectiveness of motion self-supervision
 Quantitative and qualitative results are reported to validate the effectiveness of different training strategies for the motion stream.
@@
 Thus, we evaluate its effectiveness via segmentation/prediction performance.
-From Table 7, the model with a self-supervised loss achieves better performance, which demonstrates that the prompted motion under self-supervision is beneficial to boosting detection performance.
+As validated in Table VII, the model with a self-supervised loss achieves better performance, which demonstrates that the prompted motion under self-supervision is beneficial to boosting detection performance.
 Fig.
@@
 In contrast, full-tuning the GMFlow part (Fig.
-5 (b), correspond those results in Table 3) is hardly generalized in new camouflaged scenarios.
+11 (b), correspond those results in Table III) is hardly generalized in new camouflaged scenarios.
+Fig.
+12 shows the loss curves of optical flow estimation (left) and camouflaged object segmentation (right), respectively, and one can see that both losses converge during the training process.
+Notably, the segmentation loss converges rapidly, while the flow loss converges at a slower rate.
+The straightforward utilization of a well pre-trained optical flow fundamental model results in an initially low flow loss.
+As the training proceeds, this flow loss enforces stable optimization of the segmentation-to-motion prompt learning, thereby guaranteeing a more effective prompt towards the motion stream.
 Figure 11: Visual comparisons employing different designs: (a) Our self-supervision (default EMIP), (b) Full-tuning, (c) GMFlow, (d) Ours w/o self-supervised loss.
-Table 7: EMIP without/with the self-supervised loss.
+Figure 12: Training loss of optical flow and segmentation in our proposed model EMIP. Both losses converge after about 80 epochs.
+TABLE VII: Quantitative comparison of our EMIP model with and without the self-supervised loss for optical flow.
 […]
@@
 w/ self-supervised 0.669 0.374 0.017 0.424 0.326
-Effectiveness of long-term modeling.
-As aforementioned in Table 1, by introducing the long-term consistency prompt to the short-term model EMIP, we have further enhanced the performance on all five metrics.
-Additionally, visual comparisons on the same clip in Fig.
-6 show that the incorporation of long-term consistency effectively harnesses temporal consistency, thereby reducing errors in short-term prediction.
-Figure 6: Visual comparisons on consecutive video frames.
-5 Conclusion
+IV-E Analyses of segmentation-to-motion prompt
+We introduce a segmentation-to-motion prompt strategy utilizing a camouflage feeder to enhance motion estimation.
+This approach refines the features that generate motion information, ensuring more precise responses in camouflaged regions while minimizing the impact of irrelevant information.
+To better elucidate the inner mechanism and interpretability of segmentation-to-motion prompt, as illustrated in Fig.
+13, we visualize several feature maps of $f^{flow}_{t}$, $\hat{f}^{flow}_{t}$, and $f^{2}_{t}$ in one frame.
+For a more intuitive presentation, we randomly select 10 consecutive feature maps along the channel dimension.
+As observed, the lower-level appearance features $f^{2}_{t}$ from the segmentation branch exhibits enhanced responses around the camouflaged regions.
+After being prompted by appearance features, motion features then exhibit an enhanced response towards camouflaged regions.
+This response aids in further distinguishing the motion features of camouflaged areas from the background pixels, thereby improving the motion estimation for camouflaged objects.
+Figure 13: Visualization of the segmentation-to-motion prompt.
+The left image represents the current frame.
+The red rectangles highlights prominent regions in feature maps or the current frame.
+The solid red lines mean the computation processes of the output optical flow using motion features of two adjacent frames ({$f_{t}^{flow}$,$f_{t-1}^{flow}$} or $\hat{f}_{t}^{flow}$,$\hat{f}_{t-1}^{flow}$}).
+IV-F Computational Efficiency
+To thoroughly assess model-related parameters and efficiency, we perform a comparative analysis in Table VIII against the previous cutting-edge model SLT-Net [10], under the same GPU configuration.
+Despite having $\sim$18M more parameters compared to SLT-Net, EMIP exhibits decent improvements in both detection accuracy and frames-per-second (FPS).
+Furthermore, the long-term variant of EMIP, i.e., EMIP, achieves such improvements with remarkably low amount of fine-tuned parameters (8.5M, less than 8% of the total model parameters), reducing the computational overhead during training.
+TABLE VIII: Comparison of model parameters and efficiency with previous SLT-Net, in terms of total parameters, finetuned parameters, and inference speed (frames-per-second, FPS).
+The best scores are highlighted in bold.
+Model Total Params Finetuned Params FPS $\mathcal{S}_{\alpha}\uparrow$ $F_{\beta}^{w}\uparrow$ $\mathcal{M}\downarrow$
+SLT-Net 82.38M 82.38M 5.5 0.637 0.304 0.027
+EMIP 100.86M 96.06M 7.8 0.669 0.374 0.017
+EMIP 109.04M 8.50M 6.2 0.675 0.381 0.015
+IV-G Generalization Ability on VSOD/VOS
+To thoroughly demonstrate the generalizability of EMIP, we conducted extended evaluations using the well-known VSOD/VOS datasets DAVIS16 [70], FBMS [71], ViSal [72] and SegV2 [73].
+The results, summarized in Table IX, include a comprehensive quantitative comparison against several recently published state-of-the-art methods for VSOD and VOS tasks.
+These methods represent a broad spectrum of current advancements in the field.
+Additionally, we present visual comparisons in Fig.
+14 to further illustrate the efficacy of our approach.
+Our model, EMIP, consistently excels in capturing fine contour details across various scenarios.
+For instance, it accurately delineates the foot of a dog in the 3rd row, the tail of a horse in the 4th row, and the wheel of a motorcycle in the 5th row.
+These detailed visual comparisons highlight our model’s ability to handle intricate object boundaries and maintain high fidelity in the segmentation process.
+The results clearly demonstrate that, although EMIP is specifically designed to tackle video camouflage scenes, it also performs exceptionally well in more general video segmentation tasks.
+This adaptability underscores the robustness and versatility of our approach, making it suitable for a wide range of applications beyond its original design scope.
+[…]
+Method DAVIS16 FBMS ViSal SegV2
+$\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$ $\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$ $\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$ $\mathcal{S}_{\alpha}\uparrow$ $\mathcal{F}\uparrow$ $\mathcal{M}\downarrow$
+SCOM[47]${}_{\text{TIP'2018}}$ 0.832 0.783 0.048 0.794 0.797 0.079 0.762 0.831 0.122 0.815 0.764 0.030
+MBNM[48]${}_{\text{ECCV'2018}}$ 0.887 0.861 0.031 0.857 0.816 0.047 0.898 0.883 0.020 0.809 0.716 0.026
+PDBM[49]${}_{\text{ECCV'2018}}$ 0.882 0.855 0.028 0.851 0.821 0.064 0.907 0.888 0.032 0.864 0.800 0.024
+SRP[50]${}_{\text{TIP'2019}}$ 0.662 0.660 0.070 0.648 0.671 0.134 - 0.752 0.092 - 0.683 0.095
+MESO[51]${}_{\text{TMM'2019}}$ 0.718 0.660 0.070 0.635 0.618 0.134 - - - - - -
+LTSI[87]${}_{\text{TIP'2019}}$ 0.876 0.850 0.034 0.805 0.799 0.087 0.922 0.909 0.027 0.827 0.862 0.028
+RSE[88]${}_{\text{TCSVT'2019}}$ 0.748 0.698 0.063 0.670 0.652 0.128 - - - - - -
+SSAV[89]${}_{\text{CVPR'2019}}$ 0.893 0.861 0.028 0.879 0.865 0.040 0.943 0.939 0.020 0.851 0.801 0.023
+RCR[52]${}_{\text{ICCV'2019}}$ 0.886 0.848 0.027 0.872 0.859 0.053 - - - - - -
+CAS[90]${}_{\text{TNNLS'2020}}$ 0.873 0.860 0.032 0.856 0.863 0.056 - - - 0.820 0.847 0.029
+PCSA[91]${}_{\text{AAAI'2020}}$ 0.902 0.880 0.022 0.868 0.837 0.040 0.946 0.940 0.017 0.865 0.810 0.025
+DFNet[92]${}_{\text{ECCV'2020}}$ 0.899 0.018 - 0.833 0.054 - 0.927 0.017 - - -
+[…]
+[…]
+UFO [93]${}_{\text{TMM'2023}}$ 0.874 0.797 0.032 0.868 0.803 0.041 0.940 0.914 0.012 0.836 0.746 0.057
+MAMNet [57]${}_{\text{TIP'2024}}$ 0.897 0.877 0.020 0.894 0.883 0.032 0.947 0.948 0.012 0.886 0.850 0.014
+EMIP (Ours) 0.908 0.902 0.016 0.891 0.887 0.032 0.950 0.950 0.012 0.891 0.862 0.013
+Figure 14: Visual comparisons of EMIP with four state-of-the-art VOS/VSOD methods. Red rectangles indicate challenging regions on which our EMIP excels.
+IV-H Failure Cases
+In the 1st scenario depicted in Fig.
+15, the camouflaged object is occluded.
+In this context, the detection outcome encompasses the occlusion object due to the significant resemblance across the camouflaged object, occlusion entity, and background.
+Another scenario is shown in the 2nd row of the figure.
+The displayed image represents a frame among initial frames of a sequence, lacking motion information and posing a challenge even for human visual perception, let alone learning models.
+For the 3rd scenario, the prominent green leaves in the foreground act as strong distractors and the camouflaged object is easily mistaken as part of the foreground foliage.
+It is important to note that these challenges are not unique to our model.
+Even state-of-the-art methods from previous research, such as PNS-Net [84] and SLT-Net [10], encounter similar difficulties, as shown in Fig.
+15.
+Consequently, addressing these specific challenges remains a crucial direction for future research in the field of VCOD.
+Figure 15: Some failure cases of EMIP and two most recent methods.
+V Conclusion
 We propose EMIP, an innovative framework for VCOD that explicitly handles motion cues through the utilization of a frozen pre-trained optical flow fundamental model.
@@
 The core idea of interaction between these two streams is orchestrated through an interactive prompting mechanism.
-Experimental results show the paradigm of interactive prompting of EMIP can enhance outputs of both streams, further achieving accurate prediction.
+Experimental results show that the paradigm of interactive prompting of EMIP can enhance the outputs of both streams, further achieving accurate prediction.
+Comprehensive ablation studies and in-depth discussions validate the key components of EMIP.
 In addition, we present an extended version of EMIP, incorporating historical features into the prompt to alleviate short-term prediction errors and enhance overall accuracy.
-Our contributions not only achieve compelling results on two VCOD benchmarks but also provide fresh insights into addressing the challenging VCOD task.
-We hope that the proposed framework will serve as a catalyst for inspiring further research in this emerging field.
+Moreover, the proposed EMIP is extended to the general video object segmentation task, consistently delivering improved performance and validating its generalizability and adaptability.
+Our contributions not only achieve compelling results on two VCOD benchmark datasets, but also provide fresh insights into addressing the challenging VCOD task.
+We hope that the proposed framework could serve as a catalyst for inspiring further research in this emerging field.
+We believe that making controllable and adjustable optimization prompts for fundamental models presents an intriguing avenue for future investigation.